OpenStudy (anonymous):

Raise the quantity in parentheses to the indicated exponent, and simplify the resulting expression. Express answers with positive exponents. (-33x^-2y^2 / 165x^-1) ^3

OpenStudy (pfenn1):

Is this the correct expression?\[\left(\frac{-33x^{-2}y^2}{165x^{-1}} \right) ^3 \]

OpenStudy (pfenn1):

Recall that \[x^{-a}=\frac{1}{x^a}\] and \[x^ax^b=x^{a+b}\]and that 33 and 165 are both divisible by 11.